OpenStudy (anonymous):

A point charge of -20uC is placed on the x axis at x=-2m. A second point charge of +40uC is placed on the x acis at x=+2m. a) Find the electric force between the two charges. b) What is the electric feild at x=6m on the x axis. c) Find a location on the x axis where the electric field is zero.

OpenStudy (anonymous):

I did part a and came up with 449.5x10^-3. I used the equation \[F=\frac{ k|q1||q2| }{ r ^{2} }\]

OpenStudy (anonymous):

\[F=k \frac{ (20)(2) }{ 4^2}\]

OpenStudy (anonymous):

What does this apply to hah?

OpenStudy (anonymous):

to a)

OpenStudy (anonymous):

OH! I did a haha. :) I'm lost on b though...

OpenStudy (anonymous):

\[\frac{ k(20)(40) }{ 4^2 }\]

OpenStudy (anonymous):

i meant to say that

OpenStudy (anonymous):

b)\[E=\frac{ kQ }{ r }\] \[Q=(q _{1}+q _{2})/2,r=4\]

OpenStudy (anonymous):

for b) you are looking for the resultant E you will have \[E =k q/ r ^{2}\] for the -20 uC r = 8m the \[E = 5k/26 \to the \left\] and for 2uC r=2, then \[E = k/2 \to the \right \]. the difference between the two E will you resultant electric field

OpenStudy (anonymous):

Thank you Fifi!!! I was lost on this part. Part c) is still giving me trouble. I know that I take a difference of two kq/r^2 equations but the r is different for each equation and I don't know what to set it to...

OpenStudy (anonymous):

 Add the fields, set equal to zero, solve for r, and remember that the x coord. will be x =-2 -r